OpenStudy (anonymous):

SHOW ALL WORK. The owner of a small firm has just purchased a personal computer, which she expects will serve her for the next 2 years. The owner has been told that she should buy maintenance plan to offset the chance of the computer malfunctioning. She is told that there is a 1% chance of incurring $400 in damage, 2% chance of incurring $200 in damage and a 10% chance of incurring $100 in damage. If she makes decisions based on the expected value, how much should she be willing to pay for this maintenance plan?

OpenStudy (paxpolaris):

400*0.0 + 200*0.02 + 100*10 = 4 + 4 + 10 = $18